# Virtual Reflections on a Dynamic 2D Eye Model Improve Spatial Reference Identification

### v3

@@ Abstract
The visible orientation of human eyes creates some transparency about people’s spatial attention and other mental states. This leads to a dual role for the eyes as a means of sensing and communication. Accordingly, artificial eye models are being explored as communication media in human-machine interaction scenarios. One challenge in the use of eye models for communication consists of resolving spatial reference ambiguities, especially for screen-based models. Here, we introduce an approach for overcoming this challenge through the introduction of reflection-like features that are contingent on artificial eye movements. We conducted a user study with 30 participants in which participants had to use spatial references provided by dynamic eye models to advance in a fast-paced group interaction task. Compared to a non-reflective eye model and a pure reflection mode, their combination in the new approach resulted in a higher identification accuracy and user experience, suggesting a synergistic benefit.
Keywords eye-model  $\cdot$ spatial referencing  $\cdot$ human-robot interaction
@@ 1 Introduction
First and foremost, the eyes are sensory organs for acquiring visual information from the environment. Sharp visual perception through the eyes relies on the projection of light from a scene at a depth of interest onto light-sensitive tissue inside the eye. To achieve such a projection selectively, the eyes have evolved mechanisms that allow us to sample a scene through various actions that alter, for instance, eye orientation, pupil diameter, and lens curvature. One byproduct of this action-guided visual perception is that aspects of cognition become observable. For example, eye movements and the direction of the pupil reveal where a person is looking and, when further following the gaze direction, what a person is currently attending to. Thus, eyes may not only be understood as sensory organs but also as communication devices [1, 2]. Indeed, in support of their informative value, eyes have been found to be a particularly strong attractor for other people’s gaze (see, e.g., [3, 4, 5]) and the exploitation of gaze direction transparency appears to even have evolved into reflexive gaze following behavior [6]. According to the cooperative eye hypothesis [1, 7], human eyes may have developed their particularly visible features - compared to other animals - precisely to increase their social utility. The social role of the eyes is also emphasized by the participation of the eyes in the expression of and influence on emotions [8, 9, 10] as well as the particular sensitivity to eye contact [11]. Due to the perceptual salience of the eyes, communication abilities, and social role, models that mimic human eyes are regularly used as communication tools, especially in virtual characters and robots designed for human-machine interaction (e.g., [12, 13, 14, 15, 16, 2]). If implemented successfully, such eye models can be exploited to capture a person’s attention intuitively, communicate directions of interest, and even indicate some mental states, making them promising tools for supporting human-machine cooperation [17, 18].
@@ 1.1 Challenges for communicative eye models
However, eye models on virtual characters and robots are often subject to constraints that may limit communication utility. Recreating human-like ocular behavior requires multiple actuators to be driven in a complex conjunction at low latency and with high accuracy (e.g.,  [19]). This means that design and construction are either complex or a compromise in terms of model performance. Artificial eyes can thus easily fail to recreate realistic eye movements with human-like latency, accuracy, and context-dependent variability, due to limitations of the available hardware or the behavior-driving models. In consequence, their movements can appear unnatural, scripted, or even disturbing [20, 21], possibly leading to an ignorance of the eyes in the long term. Virtual eyes that are presented on displays [14, 16, 22] may overcome many limitations related to latency and available degrees of freedom but can have difficulties in unambiguously conveying spatial information due to their two-dimensional screen projection.
@@
The ability to follow the gaze of another person is partially dependent on the spherical shape of the eyes, which creates depth cues that aid estimations about each eye’s orientation. Considering the individual orientations of both eyes relative to each other can reveal not just the direction but the actual distance and even point of attention in space because the point of convergence of both eyes’ orientation (in extension) corresponds to the focus point [23]. Because the angle of convergence becomes smaller with distance and, conversely, increases with proximity, it is a useful indicator of focus distance.
When another person is located further in the distance (thereby reducing binocular disparity) or when we only see a two-dimensional projection of another person, for example, on pictures and screens, an accurate estimation of that person’s eye orientation and focus direction is aggravated. In these cases, other contextual information, such as head orientation and body posture in relation to the environment can be used to infer another person’s region of attention [4]. However, an absence of depth information and, in consequence, also the distance cues from binocular convergence can lead to misperceptions of gaze direction and distance, especially in ambiguous situations with multiple possible regions of interest such as during group interactions. There exist some possibilities to simulate depth on two-dimensional displays by varying the image in response to changes in the location of the observer [24]. Such a simulation requires head-tracking of the observer and translation into corresponding image changes. However, the simulation of depth occurs only during observer movement and does not engage stereopsis. Furthermore, the dependence of a 3D effect on the perspective of the observer becomes problematic during interactions with groups of people, each of which occupies a different location and perspective. Unless each perspective is rendered and displayed individually, this limits the utility of perspective-adapted depth cuing to 1 to 1 human-robot interaction scenarios.
To overcome the loss of depth information while preserving the advantages of virtual eyes and other three-dimensional facial features, retro-projected robot heads have been introduced [25, 26]. These project images onto the inside of a translucent head model to realize a three-dimensional head-shaped display on which many spatial features can be conveyed.
However, such approaches can suffer from masking by ambient illumination [25]. Ambient light sources may not only make projected elements more difficult to see but the interactions of light and shadow with the head can lead to seemingly unnatural effects such as an inversion of shadows, in which regions not illuminated by outside light sources are more clearly visible than those that are. Retro-projected head and eye models may hence be restricted to environments with sufficiently controlled illumination.
@@
The discussed approaches for creating eye models for communication purposes all come with issues that constrain their utility in various ways. The following section aims to introduce an approach that may circumvent some of these issues and extend the social utility of artificial eye models.
1.2 Benefits of a smooth surface
In interactions between humans, looking at another person’s eyes is just one of two steps in trying to understand their focus of attention. At least for spatial references, the other step consists of mapping the gaze onto its target in space. The problem of inaccurate ocular cues lies in the uncertainty they create for the inferred gaze vector that would need to be traced. However, conveying the focus of attention through the eyes does not necessarily have to depend on such a gaze vector for artificial eyes. The appearance of virtual eyes represented on a display can be altered arbitrarily to support an observer’s understanding of the focus of attention. A question that arises is, how this appearance may change or what may be displayed to convey references intuitively and without interfering with other communication features.
Human eyes have a smooth outer surface that reflects a portion of the light in the scene. In close proximity and with suitable illumination an eye can appear as a partial convex mirror of the scene it is facing.
@@ 1.2 Benefits of a smooth surface
Seeing a reflection in someone’s eyes is thus a natural phenomenon, although the size of the reflection typically makes recognition of the reflected scene difficult. Artificial eyes are not restricted to the same size and reflection properties as human eyes. Reflections on such eyes may be more recognizable, especially because the use of virtual eye models also allows for virtual reflections and projections that can be altered to enhance visibility. Accordingly, here it is proposed to augment virtual eye models intended for human-machine interaction, such as eyes shown on screens or projections, by including virtual mirror images of the machine’s current focus of attention in the eye models. In reference to that feature, here we refer to such eye models as mirror eyes.
Mirror eyes simultaneously display the direction and the content of focus. For an observer, this can turn the described two-step process of understanding spatial references into a single step and remove the need for unambiguously interpretable gaze vectors, substantially facilitating the design of effectively communicating eyes. Depending on user preferences, application context and the size and visibility of a machine’s eye models, aspects of virtual mirror images, such as scene crop, size, opacity, location, and distortion may be varied.
@@ 2.1 Prototype
We created a functional prototype that implements the mirror eye concept for utilization on a wide range of ubiquitous hardware, such as personal computers with webcams. It relies on the following sub-modules:
2.1.1 Dynamic eye model
@@
The first module consists of a dynamic two-dimensional eye model that allows real-time movement of pupil and iris images relative to the sclera to create an illusion of rotation. To align these movements with locations of interest in space, a mapping from a camera image space to eye coordinates is applied. The camera image is mapped from the perspective of the location of the eye model on a screen. To simulate sensitivity to depth despite the flat camera image, we have created a module that estimates distances to objects in the scene based on provided object sizes and camera specifications for focal length and sensor size. The depth estimates are then used to adjust horizontal offsets in the case of the simultaneous application of multiple eye models. In accordance with typical vertebrate anatomy, we apply two eye models in conjunction and utilize the depth estimates to create individual horizontal offsets for each eye’s pupil and iris that create an effect of binocular disparity. This allows an observer to infer depth information from alignment differences between the eyes and may thus facilitate gaze target recognition.
2.1.2 Attention
@@
To make the eye modules come to life by moving across a scene in a meaningful way, an attention module is required. Simply put, the task of such a module is to define and continuously update target coordinates in camera space that are then mapped to pupil and iris position as described above. Because we were interested in the application of eye models as referencing tools in social settings, we chose to let the eyes look at and follow people in a scene. For this purpose, we implemented a face-tracking module that can identify and simultaneously track multiple faces present in a camera feed. Target coordinates are then defined relative to one or multiple of these faces, allowing the eyes to appear to look at, follow, and switch between faces.
2.1.3 Virtual reflection
@@
Having established dynamic eye models that appear to look at and follow faces in a scene, the next module adds the virtual reflection capability consisting of the following steps: For each eye, an overlay of a horizontally flipped camera feed at a chosen transparency level is created. To indicate what part of the scene is being attended, the pupil should be located at or near the region of interest on the respective camera image. However, only moving the full eye model across the large image space to point to objects in the scene would result in a loss of the previously described eye-scene alignment. To overcome this issue, we apply a spatial shift and coordinate translation that moves the modified camera image opposite to the direction of eye movement as depicted in Figure 2 and described by Equations 1-4 as follows.
@@
2.1.4 Software
We implemented all modules using the Python programming language [27] with the libraries OpenCV [28] for image processing and manipulation and MediaPipe [29] for face detection.
2.2 Investigation
We were interested in the utility of mirror eyes for communicating spatial references to people in a group setting.
In particular, we wanted to test the following hypotheses within that scope:
H1:
Augmenting eye models with pupil-centered virtual reflections of attended regions of interest (i.e., using mirror eyes) improves the identification accuracy of such references by people compared to eye models that lack such augmentation.
H2:
The use of mirror eyes reduces the speed of positive spatial reference identification.
H3:
The use of mirror eyes as a spatial reference tool in a group setting improves the user experience compared to eye models that lack either mirror-like augmentation or pupil-like features.
We designed an experiment that required participants within a group to quickly and accurately identify who in the group was being attended by an eye model to proceed in a fast-paced game-like task. In each experiment, we recorded gaze reference identification accuracy and speed. The purpose of the secondary game-like task was to create conditions that would allow a quick and structured recording of many samples while distracting the participants cognitively from the primary stimulus-response task.
2.2.1 Participant task
@@ 6. Target succession rule:
The introduction of a third player and the machine mediator, instead of a turn-taking rule, ensured that participants were always uncertain about whose turn it would be next and had to make a decision based on the mediator’s gaze to proceed with the game. The use of the button press was two-fold: It facilitated identifying and recording whether a participant had understood the gaze reference correctly and allowed us to measure how quickly true positive and false positive reference identifications took place. We regarded the absence of a button press within the 3-second time limit as a negative response. The short time limit for the button press urged participants to react as quickly as possible and thus reduced the potential influence by confounding factors. We defined each sequence of gazing, button press, and word utterance as one trial. Thus, a single trial took a maximum of 8 seconds.
2.2.2 Experiment structure
@@
We utilized three different display conditions of the eye models. The first condition (eye-only) consisted of pure eye models as described in Section 2.1.1 without any reflection-like properties. The second condition (mirror-only) consisted of a pure dynamic mirroring mode that was dynamically centered on the current face of interest as described in Section 2.1.3 but contained no typical eye features such as pupil and iris. The third condition (mirror-eye) consisted of the combination of the first two conditions, i.e., the overlay of dynamically updated mirror-like features on the pupil and iris. The use of these three conditions allowed us to test whether any potential effects of the mirror eyes might be attributed to one of its two major constituents or whether these would arise only from their combination.
2.2.3 Procedure
@@
Figure 3 illustrates the procedure and individual components of the experiment. Before the start of the experiment, all participants received detailed instructions and practiced for 10 trials in each eye condition.
@@
The subsequent experiment consisted of two parts that varied in continuity of exposure to each of the three conditions as well as in the total number of trials. In the first part, the conditions were kept constant for 3 blocks of 30 trials. The order of blocks was pseudo-randomized between participant triplets. In the second part, the conditions switched pseudo-randomly between individual trials, yielding at least 15 trials per condition within a single block. While blocks in the first part allowed participants to anticipate the nature of the cue they would receive from the mediator, the mixing of conditions in the second part added cue uncertainty, making tuning to a specific set of features disadvantageous. Differences between the two parts could thus point towards a possible dependence or independence between identification performance and cue type predictability.
@@
Because participants could steal individual trials from each other by wrongly pressing the button when another participant was attended, we created a real-time trial balancing procedure that would either swap or append trials for involved participants in such cases depending on compatibility with the target succession rule (see Section 2.2.1). False responses by participants could therefore extend the total number of trials in a block up to a predefined total time limit of 4 minutes for the first part and 6 minutes for the second part.
@@
2.2.4 Participants

### v4

@@ Abstract
The visible orientation of human eyes creates some transparency about people’s spatial attention and other mental states. This leads to a dual role of the eyes as a means of sensing and communication. Accordingly, artificial eye models are being explored as communication media in human-machine interaction scenarios. One challenge in the use of eye models for communication consists of resolving spatial reference ambiguities, especially for screen-based models. To address this challenge, we introduce an approach that incorporates reflection-like features that are contingent on the movements of artificial eyes. We conducted a user study with 30 participants in which participants had to use spatial references provided by dynamic eye models to advance in a fast-paced group interaction task. Compared to a non-reflective eye model and a pure reflection mode, the superimposition of screen-based eyes with gaze-contingent virtual reflections resulted in a higher identification accuracy and user experience, suggesting a synergistic benefit.
@@ 1 Introduction
First and foremost, the eyes are sensory organs that acquire visual information from the environment. Sharp visual perception through the eyes relies on the projection of light from a scene at a depth of interest onto light-sensitive tissue inside the eye. To achieve such a projection selectively, the eyes have evolved mechanisms that allow us to sample a scene through various actions that alter, for instance, eye orientation, pupil diameter, and lens curvature. One byproduct of this action-guided visual perception is that aspects of cognition become observable. For example, eye movements and the direction of the pupil reveal where a person is looking and, when further following the gaze direction, what a person is currently attending to. Thus, eyes may be understood not only as sensory organs but also as communication devices [1, 2]. Indeed, in support of their informative value, eyes have been found to be a particularly strong attractor for other people’s gaze (see, e.g., [3, 4, 5]) and the exploitation of gaze direction transparency appears to even have evolved into reflexive gaze following behavior [6]. According to the cooperative eye hypothesis [1, 7], human eyes may have developed their salient features precisely to increase their social utility. The social role of the eyes is also emphasized by the participation of the eyes in the expression of and influence on emotions [8, 9, 10] as well as the particular sensitivity to eye contact [11]. Due to the perceptual salience of the eyes, communication abilities, and social role, models that mimic human eyes are regularly used as communication tools, especially in virtual characters and robots designed for human-machine interaction (e.g., [12, 13, 14, 15, 16, 17, 2]). If implemented successfully, robotic eye models have been found to elicit similar psychophysiological responses in humans as the eyes of human interaction partners[18, 19]. Such findings suggest that eye models may be purposefully used in interactive and especially in cooperative [20, 21, 22] settings, e.g., to capture a person’s attention intuitively, communicate directions of interest, or indicate specific mental states.
@@ 1.1 Challenges for communicative eye models
However, eye models on virtual characters and robots are often subject to constraints that may limit communication utility. Recreating human-like ocular behavior requires multiple actuators to be driven in a complex conjunction at low latency and with high accuracy (e.g.,  [23]). This means that design and construction are either complex or a compromise in terms of model performance. Artificial eyes can therefore easily fail to recreate realistic eye movements with human-like latency, accuracy, and context-dependent variability, due to limitations of the available hardware or the behavior-driving models[15]. In consequence, their movements can appear unnatural, scripted, or even disturbing [24], possibly leading to an ignorance of the eyes in the long term. Virtual eyes that are presented on displays (e.g., [16, 17, 25, 26, 15]) may overcome many limitations related to latency and available degrees of freedom but can have difficulties in unambiguously conveying spatial information due to their two-dimensional screen projection.
@@
The ability to follow the gaze of another person is partially dependent on the spherical shape of the eyes, which produces visual cues that help estimate the orientation of each eye. Considering the relative orientations of both eyes can reveal not just the direction but the actual distance and even the point of attention in space because the point of convergence of both eyes’ orientation (in extension) corresponds to the focus point [27]. Because the angle at the point of convergence becomes smaller with distance and, conversely, increases with proximity, this eye vergence angle (EVA) is a useful indicator of focus distance. The EVA is closely tied to the interpupillary distance (IPD), which varies slightly as a function of the EVA and thus encodes some depth information. In natural 3D settings, this information is reinforced by motion parallax, as small head movements produce dynamic changes in the observer’s view that further disambiguate the target of another person’s gaze.
Estimating another person’s eye orientation and focus direction becomes more challenging when the eyes are distant or presented in a 2D format, such as on screens. In these cases, contextual cues, such as head orientation and body posture, can support inferences about another person’s attention region [4]. However, the lack of depth information can lead to misperceptions of gaze direction and distance, particularly in ambiguous situations like group interactions. While depth may be simulated on 2D displays through artificial motion parallax, e.g., via head-tracking contingent image changes [28, 29], such approaches are limited as they rely on observer movement and do not engage stereopsis. Moreover, adapting depth cues to individual perspectives becomes problematic in multi-person interactions, limiting its utility to one-to-one interaction scenarios.
To overcome the loss of depth information while preserving the benefits of virtual eyes and 3D facial features, researchers have developed retro-projected robot heads [30]. These project images onto a translucent head model, creating a 3D display that conveys spatial information. However, this approach is susceptible to ambient illumination [30], which can cause projected elements to be obscured and create unnatural lighting effects, such as inverted shadows. Retro-projected head and eye models may hence be restricted to environments with sufficiently controlled illumination.
@@
All discussed approaches for creating eye models for communication purposes come with issues that constrain their utility in various ways. The following section aims to introduce an approach that may circumvent some of these issues and extend the social utility of artificial eye models.
1.2 One-step spatial references
In human interactions, understanding another person’s focus of attention involves two steps: looking at their eyes and mapping their gaze onto its target in space. Uncertainty in ocular cues can make this spatial inference unreliable. However, artificial or virtual eyes are not bound by biological constraints but can be visually modified to reduce communicative uncertainty. This raises the question of how their appearance might be adapted to intuitively convey spatial references without interfering with other communicative functions.
@@ 1.2 One-step spatial references
Human eyes partially reflect their surroundings like convex mirror, although the small size of the reflection makes it difficult to recognize. In contrast, artificial eyes are not limited by size or physical reflection properties. Reflections on such eyes may be more recognizable, especially because the use of rendered eye models also allows for virtual reflections and projections that can be altered to enhance visibility. Consequently, here we propose to augment artificial eye models intended for human-machine interaction, such as eyes shown on screens, by including virtual mirror images of the machine’s current focus of attention in the eye models (see Figure 1). In reference to that feature, we refer to such eye models as Mirror Eyes.
Mirror Eyes simultaneously display the direction and content of attention. For an observer, this can turn the described two-step process of understanding spatial references into a single step and remove the need for unambiguously interpretable gaze vectors, substantially facilitating the design of effectively communicating eyes. Although the added detail displayed on such Mirror Eyes could increase information processing demands for an observer, we expect the perception of virtual reflections to be accompanied by a gain in gaze target identification accuracy, especially in situations with potential reference ambiguity, such as group interactions. We further expect the experience of such references to be enhanced in comparison to classical eye models and to pure dynamic virtual reflections that lack eye features, such as pupil and iris, which typically indicate spatial attention.
To formalize the preceding claims, we propose the following research hypotheses. The use of Mirror Eyes …
H1:
…   improves people’s identification accuracy of spatial references compared to eye models that lack mirror augmentation.
H2:
… reduces the speed of positive spatial reference identification compared to eye models that lack mirror augmentation.
H3:
…  as a spatial reference tool in a group setting improves the user experience compared to eye models that lack either mirror-like augmentation or pupil-like features.
To evaluate these hypotheses, we developed a functional prototype of an eye model with Mirror Eye capabilities and conducted a user study using this model. The remainder of this paper presents both the model and the empirical investigation in detail.
@@ 2 Methodology - Prototype
We created a functional prototype that implements the Mirror Eye concept for utilization on a wide range of hardware, such as personal computers with webcams. It relies on the following interconnected modules as depicted in Figure 2:
2.1 Dynamic eye model
@@
The first module consists of a dynamic 2D eye model that allows real-time movement of pupil and iris images relative to the sclera to create an illusion of rotation. To align these movements with locations of interest in space such that the pupil appears to point towards a target location, a mapping from a camera image space to eye coordinates is applied (see Figure 3 and Equations 3-4). The camera image is mapped from the perspective of the location of the eye model on a screen. To simulate sensitivity to depth despite the flat camera image, we have created a module that estimates distances to objects in the scene based on provided object sizes and camera specifications for focal length and sensor size. The depth estimates are then used to adjust horizontal offsets in the case of the simultaneous application of multiple eye models. In accordance with typical vertebrate anatomy, we apply two eye models in conjunction and utilize the depth estimates to create individual horizontal offsets for each eye’s pupil and iris such that interpupillary distance varies with focus distance, resulting in an effect of binocular disparity. This allows an observer to infer depth information from alignment differences between the eyes and thus may facilitate gaze target recognition.
Figure 2: Connected modules of the Mirror Eye prototype.
2.2 Attention
@@
To make the eye modules come to life by moving across a scene in a meaningful way, an attention module is required. Simply put, the task of such a module is to define and continuously update target coordinates in camera space that are then mapped to pupil and iris position as described above. Because we were interested in the application of eye models as referencing tools in social settings, we chose to let the eyes look at and follow people in a scene. For this purpose, we implemented a face-tracking module that can detect and simultaneously track multiple faces present in a camera feed. For the purpose of the present work, we based the face-tracking on the spatiotemporal proximity between individual face detections. With this method, faces detected at subsequent time steps near previous detections in image space are assigned the same identifier. This enables smooth tracking without the need for individual face learning or recognition for scenarios in which the minimal distance between multiple people can be well controlled. Target coordinates are then defined relative to one or multiple recent coordinates of face identifiers, allowing the eyes to look at, avoid, follow, and switch between faces.
2.3 Virtual reflection
@@
Having established dynamic eye models that appear to look at and follow faces in a scene, the next module adds the virtual reflection capability consisting of the following steps: For each eye, an overlay of a horizontally flipped camera feed is created at a chosen transparency level. To indicate what part of the scene is being attended, the pupil should be located at or near the region of interest on the respective camera image. However, only moving the full eye model across the large image space to point to objects in the scene would result in a loss of the previously described eye-scene alignment. To overcome this issue, we apply a spatial shift and coordinate translation that moves the modified camera image opposite to the direction of eye movement as depicted in Figure 3 and described by Equations 1-4 as follows.
@@
2.4 Software
We implemented all modules using the Python programming language [31] with the libraries OpenCV [32] for image processing and MediaPipe [33] for face detection.
3 Methodology - Experiment
To test the hypotheses specified in Section  1.2, we designed an experiment that required participants within a group to quickly and accurately identify who in the group was being attended by an eye model to proceed in a fast-paced game-like task. In each experiment, we recorded the accuracy and speed of gaze reference identification. The purpose of the secondary game-like task was to create conditions that would allow a quick and structured recording of many samples while distracting the participants cognitively from the primary stimulus-response task.
3.1 Participant task
@@ 6. Target succession rule:
The introduction of a third player and the machine mediator, instead of a turn-taking rule, ensured that participants were always uncertain about whose turn it would be next and had to make a decision based on the mediator’s gaze to proceed with the game. The use of the button press was two-fold: It facilitated identifying and recording whether a participant had understood the gaze reference correctly and allowed us to measure how quickly true positive and false positive reference identifications took place. We considered the absence of a button press within the 3-second time limit as a negative response. The short time limit for the button press urged participants to react as quickly as possible, thereby reducing the potential influence of confounding factors and enabling rapid data collection. We defined each sequence of gazing, button press, and word utterance as one trial. Thus, a single trial took a maximum of 8 seconds.
3.2 Experiment structure
@@
We utilized three different display conditions of the eye models as depicted in Figure 4. The first condition (Eyes-Only) consisted of pure eye models as described in Section 2.1 without any reflection-like properties. The second condition (Mirrors-Only) consisted of a pure dynamic mirroring mode that was centered on the current face of interest as described in Section 2.3 but contained no typical eye features such as pupil and iris. The third condition (Mirror Eyes) consisted of the combination of the first two conditions, i.e., the overlay of dynamically updated mirror-like features on the pupil and iris. The use of these three conditions allowed us to test whether any potential effects of the Mirror Eyes might be attributed to one of its two major constituents or whether these would arise only from their combination.
(a) Eyes-Only
(b) Mirrors-Only
(c) Mirror Eyes
Figure 4: Eye model display conditions. 4a: Basic eye model with no reflection of an attended face. 4b: dynamic mirror that tracks an attended face. 4c: Combination of eye model and dynamic mirror of attended face on pupil and iris (Mirror Eyes). The faces in Figures 4b and 4c have been blurred for publication purposes to protect privacy. No face blur was present in the experiment.
3.3 Procedure
@@
Figure 5 illustrates the procedure and individual components of the experiment. Before starting the experiment, all participants received detailed instructions and practiced for 10 trials in each eye condition.
@@
The subsequent experiment consisted of two parts that differed in how trials of the different eye model display conditions were grouped into blocks. In the first part, each display condition was used for a block of 30 trials, resulting in three blocks. Blocks of this type are further referred to as ”single condition block”. The order of single condition blocks within the first part of the experiment was pseudo-randomized between participant triplets. The second part of the experiment consisted of a single ”mixed condition block”. In a mixed condition block the eye model display conditions switched pseudo-randomly between individual trials, such that the eyes might, for instance, look at one person with Mirror Eyes but focus on the next with non-reflective eyes. At least 15 trials of each condition were distributed across a mixed condition block. While the single condition blocks in the first part allowed participants to anticipate the nature of the cue they would receive from the mediator, the mixing of conditions in the second part added cue uncertainty, making tuning to a specific set of features disadvantageous. Differences between the two parts could thus point towards a possible dependence or independence between identification performance and cue type predictability.
@@
Because participants could steal individual trials from each other by wrongly pressing the button when another participant was attended, we created a real-time trial balancing procedure that would either swap or append trials for involved participants in such cases depending on compatibility with the target succession rule (see Section 3.1). False responses by participants could therefore extend the total number of trials in a block up to a predefined total time limit of 4 minutes for the first part and 6 minutes for the second part.
Assuming a medium effect size of $\eta^{2}=0.06$, a power analysis for a repeated measures design with three within-subject conditions, 30 participants, and an alpha level of 0.05 estimated the statistical power to be 84.9%.
@@
(a) Experiment stimuli presented on a TV display
(b) Schematic of participants’ positions
Figure 6: Schematic of the experimental setup: 6a: Arrangement and dimensions of the visual stimuli consisting of one of the three dynamic eye models rendered into the eyes of a mock-up robot head shown on a gray display. 6b: Distancing between the three participants. The participants were free to move their heads and bodies while in their positions.
3.4 Dependent measures
3.4.1 Reference identification performance
To test H1 and H2, we used the reference identification accuracy and reaction times as dependent measures. Both are derived from the participants’ button press behavior. The reference identification accuracy for each participant is defined as binary classification accuracy, i.e., as the ratio of the sum of true positive ($TP$) and true negative ($TN$) identifications over the sum of all trials involving that participant.
$\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$ (5)
Therefore, the accuracy does not only take the true positive identification of a reference into account but is also affected by the participants’ ability to identify when they are not being attended. To provide opportunities for the identification of such non-references, we programmed the eye model to ”make mistakes” in a subset of trials by not attending to any particular participant but instead focusing on areas located between participants. For such trials, the lack of a button press within the time limit would then count as a true negative identification. To ensure exposure of each participant to at least one error case for each block and condition while minimizing error predictability, the rate of machine mistakes was set to 10% (3 trials) for each condition block in experiment part 1 and to 20% (3x3 trials) for part 2, which contained all three conditions.
We chose to use only a single response button per participant to avoid any coordination-dependent confounders. This button indicated a positive response. Negative responses were given implicitly by not responding within the 3-second time limit. Therefore, reaction times were only available for true positive and false positive responses as indicated by Hypothesis 2.
3.5 User experience
In this experiment, the 8-item version of the User Experience Questionnaire (UEQ-S; [34]) was used to assess participant experience with the virtual eyes. After completing all experimental blocks, participants filled out the UEQ-S, which consists of eight items, four representing a pragmatic scale and the other four representing a hedonic scale. Each item is rated on a 7-point semantic differential scale, ranging from -3 (most negative evaluation) to +3 (most positive evaluation). Among psychometrically validated tools for measuring user experience (e.g., [35, 36, 34]), we selected the UEQ-S because it efficiently captures two core user experience dimensions, the pragmatic and hedonic quality [37], while minimizing participant burden through its brief completion time.
3.6 Participants